Two long bar magnets are placed with their axes coinciding in such a way that the north pole of the first magnet is 2.0 cm from the south pole of the second. If both the magnets have a pole strength of 10 A-m, find the force exerted by one magnet on the other.

Grade:upto college level

1 Answers

Navjyot Kalra
9 years ago
Sol. m base 1 =m base 2 = 10 A-m r = 2 cm = 0.02 m we know force exerted by tow magnetic poles on each other = μ base 0/4π = m base 1m base 2/r^2 = 4π * 10^-7 *10^2/4π * 4 * 10^-4 = 2.5 * 10^-2 N